### v3

# CCaaLF: Concurrency Control as a Learnable Function

Abstract.
Concurrency control (CC) algorithms are important in modern transactional databases, as they enable high performance by executing transactions concurrently while ensuring correctness. However, state-of-the-art CC algorithms struggle to perform well across diverse workloads, and most do not consider workload drifts.
In this paper, we propose CCaaLF (Concurrency Control as a Learnable Function), a novel learned concurrency control algorithm designed to achieve high performance across varying workloads. The algorithm is quick to optimize, making it robust against dynamic workloads. CCaaLF learns an agent function that captures a large number of design choices from existing CC algorithms. The function is implemented as an efficient in-database lookup table that maps database states to concurrency control actions. The learning process is based on a combination of Bayesian optimization and a novel graph reduction algorithm, which converges quickly to a function that achieves high transaction throughput. We compare CCaaLF against five state-of-the-art CC algorithms and show that our algorithm consistently outperforms them in terms of transaction throughput and optimization time.
1. Introduction
Concurrency control (CC) algorithms allow database transactions to exploit hardware concurrency while ensuring correctness, i.e., transaction isolation. As modern database systems are no longer disk-bounded, CC is essential for achieving high transaction processing performance. Popular CC algorithms, such as two-phase locking (2PL) and optimistic concurrency control (OCC), do not perform well across all workloads (DBLP:journals/pvldb/YuBPDS14/1000_core_cc, ; DBLP:conf/cidr/TangJE17/ACC, ). In particular, 2PL generally works better than OCC in high-contention workloads, but the reverse is true in low-contention workloads.
Recent works on CC focus on combining multiple classical CC algorithms to support different workloads. The common approach follows a classify-then-assign process, in which transaction operations are classified into different types, then the most appropriate CC algorithm is selected for each type. The classification can be based on the transaction (DBLP:conf/mascots/TaiM96/AEP, ; DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ; DBLP:conf/sosp/XieSLAK015/Callas, ), or based on the data (DBLP:conf/sigmod/ShangLYZC16/HSync, ; DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ). The selection of CC algorithm is either heuristic (DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/mascots/TaiM96/AEP, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ) or based on learning (DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ). We note that existing works do not perform well for dynamic workloads that can drift over time. In particular, these works require manual, time-consuming changes to the classification and algorithm selection when there are workload drifts. While CormCC (DBLP:conf/usenix/TangE18/CormCC, ) and ACC (DBLP:conf/cidr/TangJE17/ACC, ) propose dynamic mechanisms that alleviate the cost of manual changes, they only support a small number of CC algorithms. Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) uses an evolutionary algorithm to learn the optimal combination of CC algorithms. It learns CC actions — e.g., to wait for dependent transactions to reach specific steps, or to expose uncommitted operations — for each data access. However, it does not support general transactions, and does not work well for dynamic workloads due to the long reoptimizing time.
Our goal is to design a learned CC algorithm that can achieve high performance across different workloads, and is efficient to optimize. More specifically, the algorithm can extract a high degree of concurrency from different workloads, while incurring small overhead. In addition, by quickly optimizing when the workload changes, it can perform well under dynamic workloads. Designing such a novel CC algorithm is challenging for three reasons. First, a large number of CC algorithms proposed in the literature are optimized for specific workloads. Thus, for our algorithm to perform well across different workloads, it must incorporate design choices from as many existing CC algorithms as possible. Second, to achieve high performance, the algorithm’s overhead must be much smaller than the transaction execution time. This is non-trivial because, in modern in-memory databases, transactions can be executed in microseconds (or even in nanoseconds). Third, learning the optimal action for each transaction operation is challenging because most of the learned parameters (such as the wait parameters in Polyjuice) are non-continuous. This lack of locality, as demonstrated in previous works (DBLP:conf/osdi/WangDWCW0021/polyjuice, ), renders machine learning techniques like Policy Gradient ineffective. As a result, costly optimization strategies, such as evolutionary algorithms and random search, are used.
We present CCaaLF, a novel learned CC algorithm that achieves our goal above. It addresses the first challenge by proposing a unified model of existing concurrency control algorithms. In particular, CCaaLF models a concurrency control algorithm as a learnable agent function :${\mathcal{{F}}}:\{{\mathbf{{s}}}\}\to\{{\mathbf{{a}}}\}$ that maps the database state $\mathbf{{s}}$ to a set of concurrency control actions $\mathbf{{a}}$. It captures existing classify-then-assign approaches: $\mathbf{{s}}$ represents the features used for classification, $\mathbf{{a}}$ indicates the actions of the assigned CC algorithm. In CCaaLF, a state $\mathbf{{s}}$ consists of features used by existing adaptive concurrency controls, such as data hotness (DBLP:conf/icde/LinCZ18/ASOCC, ) and number of dependent transactions in dependency graph (DBLP:journals/pvldb/TianHMS18/LDSF, ). An action $\mathbf{{a}}$ consists of a combination of conflict detection and resolution mechanisms. Instead of selecting from existing algorithms, we propose deconstructing concurrency control into a sequence of conflict detection and resolution actions, which allows for a broad range of CC designs by enabling arbitrary combinations of these actions. Our model is more general than existing adaptive concurrency control algorithms, such as Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) and bLSDF (DBLP:journals/pvldb/TianHMS18/LDSF, ). The former only maps data access IDs to pipeline-related wait actions, and the latter only maps dependency set sizes to priority-related wait actions. Figure 1 illustrates how Polyjuice treats all transactions accessing the same data as equals, and consequently takes a suboptimal action that results in high abort rates. In this example, bLSDF would give higher priority to transactions with more dependent transactions. However, it does not consider transaction pipelining which enables efficient transaction execution.
Figure 1. An example of CC designs not covered in Polyjuice.
CCaaLF addresses the second challenge by optimizing the cost of collecting the state $\mathbf{{s}}$ and of executing $\mathcal{{F}}$($\mathbf{{s}}$). For the former, it avoids states that are expensive to compute, e.g., hash values (DBLP:conf/icde/Su0Z21/C3, ) and conflict graphs (DBLP:conf/sigmod/WangMCYCL16/ic3, ). Instead, it uses states that can be computed in a lock-free manner, such as the number of direct dependencies and the number of executed operations. CCaaLF implements the agent function as an in-database lookup table recording the state-action pairs. This table only consumes a small amount of memory, as CCaaLF learns an efficient feature selector that reduces the input states to only hundreds of values. The whole table indeed fits in the database cache, and CC actions for an operation can be determined by a single table lookup, which only costs hundreds of CPU cycles.
CCaaLF addresses the third challenge in three steps. First, it reduces the cost of evaluating the agent function. Previous works evaluate the function by loading the lookup table into the database and monitoring the system performance. This process takes seconds to minutes (DBLP:conf/osdi/WangDWCW0021/polyjuice, ). CCaaLF adopts a Bayesian optimization approach that trains a surrogate model for predicting the system performance given different action lookup tables. By sampling from this surrogate model, it can quickly explore the search space and guide the optimization process on which agent function to evaluate next. Second, it re-formulates the original learning task of pipeline-wait actions as an optimal conflict graph learning task, which exhibits locality and therefore is amenable to common learning techniques. This is important because directly learning the pipeline-wait actions is difficult due to the lack of locality around the optimal combination of them. In particular, a small modification to an action could lead to dependency cycles and significant performance reduction. CCaaLF learns the table indirectly by solving a conflict graph learning problem, for which we design an efficient algorithm similar to genetic search. Third, CCaaLF speeds up the learning process by prioritizing the actions that yield higher improvements. Specifically, we prioritize fine-tuning wait-related actions that can bring higher performance improvements over actions like timeouts and back-off durations.
[…]
In summary, we make the following contributions:
  • We present CCaaLF, a novel learned concurrency control algorithm that achieves high performance across diverse workloads. The algorithm is also quick to optimize, which enables it to adapt quickly to workload drifts.
  • We propose a novel model of concurrency control, which is a function mapping database states into conflict-handling actions. This model captures a large number of existing CC algorithms.
  • We design an efficient learning algorithm that quickly finds a high-performance concurrency control model.
  • We conduct extensive evaluation of CCaaLF, comparing it against five state-of-the-art CC algorithms. Experimental results demonstrate that CCaaLF consistently outperforms the baselines both in throughput and in optimization time.
The remainder of the paper is structured as follows. Section 2 discusses the design space of concurrency control algorithms. Section 3 provides an overview of CCaaLF. Section 4 details how CCaaLF performs fine-grained concurrency control via an optimized action lookup table. Section 5 describes the learning process. Section 6 presents the experimental evaluation. Section 7 discusses related works, and Section 8 concludes.
2. Concurrency Control Design Space
In this section, we first describe the design space of concurrency control (CC) algorithms. We discuss how existing algorithms fit into the space, and identify the gaps that represent new opportunities.
A transaction is modeled as a sequence of data access operations, each being either a read or a write. CC algorithms allow transactions to be executed concurrently and correctly (DBLP:conf/sigmod/AdyaGLM95/occ, ; DBLP:conf/sosp/TuZKLM13/silo, ; DBLP:reference/db/Fekete18c/ssi, ; DBLP:conf/sigmod/LimKA17/cicada, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:journals/pvldb/WangK16/mocc, ; DBLP:conf/sigmod/GuoWYY21/bamboo, ; DBLP:conf/eurosys/MuAS19/DRP, ; DBLP:journals/pvldb/FaleiroAH17/EWV, ; 2PL_no_wait, ; DBLP:journals/tods/RosenkrantzSL78/2PL_wait_die, ; DBLP:journals/pvldb/BarthelsMTAH19/2PL_bound_wait, ; DBLP:journals/tods/HsuZ92/cautious_wait, ; DBLP:conf/icde/FranaszekRT91/wait_depth_limit, ; DBLP:conf/sigmod/ShengTZP19/abort_prediction, ; DBLP:conf/pods/CareyKL90/half_half, ; DBLP:conf/vldb/MoenkebergW92/opt_DMP, ; DBLP:conf/icde/MoenkebergW91/conflict_ratio, ; DBLP:conf/sigmod/ThomsonDWRSA12/Calvin, ; DBLP:journals/pvldb/KallmanKNPRZJMSZHA08/HStore, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/sosp/XieSLAK015/Callas, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:journals/pvldb/LuYCM20/Aria, ; DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/sigmod/ShangLYZC16/HSync, ; DBLP:conf/icde/Su0Z21/C3, ; NeurDB, ). To ensure correctness, a CC algorithm employs CC actions — algorithmic steps that ensure safe interleaving of data access operations among concurrent transactions. In the following, we describe a design space of CC actions consisting of two design dimensions: conflict detection and conflict resolution. A CC algorithm can use one or multiple actions drawn from this space.
(a) No conflict detection
(b) Detect critical conflicts (abort)
(c) Detect critical conflicts (block)
(d) Detect all conflicts
Figure 2. Examples of different conflict detection approaches. $T_{2}$ has RW conflicts with $T_{1}$ on A and B.
2.1. Conflicts Detection
A conflict occurs when two operations from two concurrent transactions access the same data object, and one of the accesses is a write. Existing CC algorithms use different approaches to detect conflicts, based on their assumptions about the distribution and probability of conflicts occurring during transaction execution. We identify three design choices for conflict detection: no detection, detecting critical conflicts, and detecting all conflicts. The first option optimistically assumes that there are no conflicts and therefore implements no conflict detection. The third option pessimistically assumes many conflicts and therefore detects them as soon as possible to avoid costly aborts. The second option is in between: it detects some critical conflicts while ignoring others.
We consider two transactions $T_{1}$ and $T_{2}$ executing concurrently. They both perform read and write on data objects A and B. Figure 2 illustrates examples where $T_{2}$ has Read-Write (RW) conflicts with $T_{1}$, and $T_{2}$ uses different conflict detection approaches.
No detection. The CC algorithm assumes conflicts are rare and performs no conflict detection during execution. It relies on validation at commit time for correctness. The optimistic concurrency control (OCC) and its variants follow this approach. In particular, each transaction keeps its writes locally (as opposed to exposing them to other transactions), and assumes no other concurrent transactions will update its read data or read its uncommitted writes. The algorithm allows transaction execution to proceed without the overhead of conflict detection but incurs overhead (in aborts) when there are conflicts. Figure 2(a) shows an example where $T_{2}$ aborts during the final validation because $T_{1}$ updated its previously read A and B.
Detect critical. The CC algorithm avoids the cost of aborts by identifying conflicts that are likely to cause aborts based on heuristic rules. They are called critical conflicts, which may lead to validation failure or to dependency cycles. Upon detection, the algorithm immediately aborts the transaction, e.g., ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ) and SSI (DBLP:reference/db/Fekete18c/ssi, ), or reorders operations to allow conflicting transactions to commit together, e.g., Bamboo (DBLP:conf/sigmod/GuoWYY21/bamboo, ) and IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, ). In Figure 2(b), $T_{2}$ detects that $T_{1}$ updated the previously read A during early validation and immediately aborts without executing to the end. In Figure 2(c), $T_{2}$ detects that the immediate read of B would form a dependency cycle between $T_{1}$ and $T_{2}$. It then reorders its read of B to occur after $T_{1}$’s write of B, thus breaking the cycle.
Detect all. The CC algorithm detects all conflicts that may lead to transaction aborts. The classical 2PL and its variants (2PL_no_wait, ; DBLP:journals/tods/RosenkrantzSL78/2PL_wait_die, ; DBLP:journals/pvldb/BarthelsMTAH19/2PL_bound_wait, ) adopt this approach. Figure 2(d) shows an example in which this approach causes $T_{2}$ to block for a long time, which increases transaction latency. Furthermore, the detection algorithm introduces some overhead because it has to find all conflicts.
2.2. Conflicts Resolution
Once a conflict is detected, the concurrency control algorithm resolves it by coordinating accesses from the conflicting transactions. We identify two design choices for conflict resolution: timeout, and priority.
Timeout. When the CC algorithm detects a conflict, it blocks the current transaction for a specific duration (a timeout), waiting for the conflicting operations to end in this duration, and aborts it when the timeout is exceeded. A timeout of 00 means that the transaction aborts immediately (2PL_no_wait, ), which assumes pessimistically that the transaction encountering the conflict will eventually abort. A large timeout, or timeout of $\infty$, means the transaction is blocked. In particular, the transaction is either put on a wait list associated with the data object, or is entering a spinning loop until the data object is available. This approach is implemented in lock-based and pipeline-based CC algorithms, e.g., 2PL and IC3. The blocking overhead can be significant, especially when the transaction eventually aborts.
Priority. When multiple transactions are blocked on the same data object, the algorithm may order them based on their priorities. When the object becomes available, the highest-priority transaction is woken up first. Most existing concurrency control algorithms consider no priority (DBLP:conf/sosp/TuZKLM13/silo, ; DBLP:journals/tods/HsuZ92/cautious_wait, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ), or simple FIFO priority (DBLP:journals/pvldb/BarthelsMTAH19/2PL-BW, ). Recent works heuristically assign priorities to conflicting transactions (DBLP:conf/eurosys/HuangMW17/VProfiler, ; DBLP:conf/sigmod/HuangMSW17/VATS, ; DBLP:journals/pvldb/TianHMS18/LDSF, ). For example, bLDSF (DBLP:journals/pvldb/TianHMS18/LDSF, ) gives higher priorities to transactions with higher numbers of dependent transactions. Other algorithms (DBLP:conf/eurosys/HuangMW17/VProfiler, ; DBLP:conf/sigmod/HuangMSW17/VATS, ) consider transaction age and the number of held locks when assigning priorities.
Table 1. Comparison of state-of-the-art concurrency control algorithms and CCaaLF in the design space. Timeout is omitted because it is supported by all algorithms.
Resolve w/o tuned priority Resolve with tuned priority
No detection Detect critical Detect all Detect critical Detect all
OCC (DBLP:conf/sigmod/AdyaGLM95/occ, ), Silo (DBLP:conf/sosp/TuZKLM13/silo, )
2PL (2PL_no_wait, ), TSO (DBLP:books/aw/BernsteinHG87/tso1, ; DBLP:conf/icde/LeisK014/tso2, ), MVCC (wu2017empirical/mvcc, )
EWV (DBLP:journals/pvldb/FaleiroAH17/EWV, ), IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, )
bLDSF (DBLP:journals/pvldb/TianHMS18/LDSF, ), VATS (DBLP:conf/sigmod/HuangMSW17/VATS, )
MCC (DBLP:conf/sosp/XieSLAK015/Callas, ), Tebaldi (DBLP:conf/sigmod/SuCDAX17/Tebaldi, )
ACC (DBLP:conf/cidr/TangJE17/ACC, ), C3 (DBLP:conf/icde/Su0Z21/C3, ), CormCC (DBLP:conf/usenix/TangE18/CormCC, )
ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ), Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, )
CCaaLF
Figure 3. CCaaLF’s concurrency control (CC) execution workflow and function optimization workflow.
2.3. State-of-the-Art Concurrency Control Algorithms
Table 1 compares state-of-the-art concurrency control algorithms within our design space. We omit timeout from the table because it is supported by all algorithms. It can be seen that some algorithms, such as the classic OCC and 2PL, support only one type of CC actions. Some recent algorithms, such as Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) and ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ), support multiple CC actions. It allows them to fine-tune conflict handling actions based on the workload and extract more concurrency from running transactions.
The table reveals several gaps in the design space, i.e., there are CC actions not covered by existing algorithms. In particular, no algorithms consider priority-based conflict resolution in combination with detecting critical conflicts. Most algorithms, including the ones that support multiple CC actions, do not consider transaction priorities. CCaaLF introduces a novel learning-based approach for concurrency control, enabling support for all CC actions in the design space.
3. CCaaLF Overview
CCaaLF is a novel learned concurrency control algorithm that achieves high performance across different workloads and is efficient to optimize. Figure 3 shows the main components of CCaaLF: a collector that tracks real-time database states, a feature selector $\mathcal{{E}}$ that translates the collected states into a state vector $\mathbf{{s}}$, an in-database cache of an agent function $\mathcal{{F}}$ in the form of an action lookup table, and an optimizer that detects workload drifts and optimizes the agent function accordingly. We list the notations in Table 2.
CCaaLF operates across two workflows: the CC execution workflow and the agent function optimization workflow. The former is on the critical path of transaction execution and is invoked before every data access operation to determine the optimal concurrency control actions. This workflow takes the current operation $op$ and transaction $T$ as input, and outputs a vector of concurrency control actions $\mathbf{{a}}$ for the transaction manager to execute. Specifically, it consists of three steps. First, the transaction manager sends $(T,op)$ to the collector, which extracts database states based on $T$ and $op$ and combines them into a feature vector $\mathbf{{x}}$. Next, the feature selector $\mathcal{{E}}$ selects important features and generates a state $\mathbf{{s}}$, i.e., ${\mathbf{{s}}}=\mathcal{{E}}(\mathbf{{x}})$. Finally, the concurrency control actions $\mathbf{{a}}$ are retrieved from the lookup table, i.e., $\mathbf{{a}}=\mathcal{{F}}(\mathbf{{s}})$. This workflow is efficient enough such that it can be executed at per-operation granularity. We describe it in more detail in Section 4.
Table 2. Symbols and descriptions
$T$ the currently executing transaction
$op$ the currently executing transaction operation
${\mathbf{{x}}}$ the vector with all features
${\mathbf{{s}}}$ the database state
${\mathbf{{a}}}$ the CC actions for the current operation
${\mathcal{{F}}}$ the agent function
${\mathcal{{E}}}$ the feature selector
$\mathcal{{F}}$) the system performance achieved using $\mathcal{{F}}$
The optimization workflow is off the critical path of transaction execution. It maintains system performance under workload drifts, by monitoring the system performance and reoptimizing the action lookup table when necessary. In this workflow, data pairs of the current function $\mathcal{{F}}$ and the system performance $\text{Score}(\mathcal{{F}})$ are collected periodically. The data is analyzed to detect performance degradation caused by workload drifts. When workload drift is detected, CCaaLF starts optimizing the agent function for the current workload. This optimization process runs with a pre-defined time budget $t_{max}$. In the end, the newly optimized $\mathcal{F}_{next}$ is loaded into the database cache as a new action lookup table. We describe this workflow in more detail in Section 5.
4. CCaaLF Execution
In this section, we describe the CC execution workflow in CCaaLF, which lies on the critical path of transaction execution. We first present a model of concurrency control as a learnable function $\mathcal{{F}}$ that maps current database states to concurrency control actions. Next, we discuss how to efficiently capture the input of $\mathcal{{F}}$ and evaluate it quickly. Finally, we present the detailed algorithm steps.
4.1. Concurrency Control As Learnable Function
Concurrency control algorithm can be modeled as a function $\mathcal{{F}}$ that maps the current database state $\mathbf{{s}}$ to a combination of conflict detection and conflict resolution actions $\mathbf{{a}}$:
(1) $\mathcal{{F}}(\mathbf{{s}})=\mathbf{{a}}.$
Before executing an operation, CCaaLF captures current database states $\mathbf{{s}}$, computes $\mathcal{{F}}$($\mathbf{{s}}$), and executes the resulting actions $\mathbf{{a}}$. After all operations are executed, CCaaLF starts a validation process before committing the transaction. This validation process detects and resolves previously ignored conflicts, and aborts the transaction when the conflicts cannot be resolved, for example, when the transaction reads a data object from an aborted transaction.
The function $\mathcal{{F}}$ covers many existing CC algorithms, including Polyjuice, ASOCC, 2PL, LSDF, OCC, and IC3. We demonstrate it with two examples of 2PL (wait die) and ASOCC. Let $T_{req}$ and $T_{owner}$ be the timestamps of the current transaction and of the oldest lock owner. Let $hotness$ denote the hotness level of the data object being accessed. 2PL and ASOCC can be modeled as follows:
$\mathcal{F}_{\text{2PL}}(T_{req},T_{owner})=\left\{\begin{aligned} (\text{{detect all}},0,\text{FIFO pri.})&&T_{req}>T_{owner}\\ (\text{{detect all}},\infty,\text{FIFO pri.})&&\text{otherwise}\\ \end{aligned}\right.$
$\mathcal{F}_{\text{ASOCC}}(hotness)=\left\{\begin{aligned} (\text{{no detection}},0,\text{No pri.})&&hotness=\text{cold}\\ (\text{{detect all}},\infty,\text{No pri.})&&hotness=\text{hot}\\ (\text{{detect critical}},0,\text{No pri.})&&\text{otherwise}\\ \end{aligned}\right.$
Each action consists of a conflict detection technique, a timeout value, and a transaction priority (the no priority option means all transactions have the same priority). In these examples, $T_{req}$, $T_{owner}$, and $hotness$ constitute the input states $\mathbf{s}$ of 2PL and ASOCC, respectively. The function $\mathcal{F}_{\text{2PL}}$ and $\mathcal{F}_{\text{ASOCC}}$ output actions which are then executed by the CC algorithm. In particular, 2PL aborts the transaction when $T_{req}$ is higher than $T_{owner}$, otherwise, it blocks the transaction and wakes it up later in a FIFO order. ASOCC ignores conflicts for cold data. It detects all conflicts and blocks the transaction for hot data. In other cases, it aborts the transaction when it finds critical conflicts.
4.2. State Capturing
Figure 4. CCaaLF’s state capturing process.
The input $\mathbf{{s}}$ of $\mathcal{{F}}$ is important as it determines the quality of CC actions, and it contributes to the CC algorithm’s overhead. There are two challenges in capturing $\mathbf{{s}}$. First, it must be done quickly since it is on the critical path of transaction execution. Modern in-memory databases can commit a transaction in microseconds, therefore $\mathbf{{s}}$ should be captured in order of nanoseconds. Second, the input must contain relevant database states so that the optimization of CC actions can extract a high degree of concurrency. Furthermore, the input space must be small to minimize storage overhead and to speed up the optimization of $\mathcal{{F}}$.
Figure 4 shows the state-capturing process in CCaaLF. It employs two optimizations to address the challenges above. The first optimization is to use a fixed set of lightweight features that can be efficiently captured across transactions in a lock-free manner. CCaaLF uses nine features, grouped into transaction features and graph features, as shown in Figure 4. These features have been shown to be effective at capturing transaction and dependency graph characteristics (DBLP:journals/pvldb/TianHMS18/LDSF, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/sigmod/HuangMSW17/VATS, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ). More importantly, they are common features readily available in transactional databases, and therefore are not specific to any concurrency control algorithm. The second optimization leverages a feature selector $\mathcal{{E}}$ that removes non-influential features. The result is a small set of states that is used as the input for $\mathcal{{F}}$. The selected features achieve a balance between the quality of the optimized agent function and the optimization time. CCaaLF provides a default feature selector that achieves good performance. The user can improve this default selector, at some cost, by running the optimization algorithm described in Section 5.5.
Figure 5. CCaaLF’s action lookup table.
4.3. Function Evaluation
The challenge in designing $\mathcal{{F}}$ is to ensure both efficiency and effectiveness. Specifically, the function must be quick to compute, since it is on the transaction’s critical path. Furthermore, the output actions must be able to extract high degrees of concurrency across different workloads. For efficiency, CCaaLF implements the function as a learnable action lookup table, such that the function computation reduces to a single table lookup. Although implementing $\mathcal{{F}}$ as complex machine learning models can approximate the state-action mapping more accurately, the inference cost is too high for concurrency control. In contrast, by learning the agent function in the form of a lookup table, the actions can be computed in only hundreds of CPU cycles.
To achieve high performance across diverse workloads, $\mathcal{{F}}$ learns combinations of design choices from existing CC algorithms optimized for different workloads. An output of $\mathcal{{F}}$ contains a conflict detection mechanism, which can be either no detection, detect critical, or detect all. It also contains a timeout value and a transaction priority for conflict resolution. Figure 5 shows an action lookup table where the first column is the input state and the other columns make up the CC actions. We will discuss the columns in more detail in Section 4.5. CCaaLF implements several optimizations for the agent function, including pruning suboptimal functions from the search space, learning transaction backoffs, and transforming some actions into more learnable forms.
Assign high priority after detect all. Unlike 2PL, in CCaaLF, it is possible for an operation that passes the detect all check to be ordered after another operation with a higher priority, leading to its eventual abort and retry even if both of them are using detect all action. Thus, we heuristically assign the highest priority to transactions that have passed the detect all check, avoiding such unnecessary aborts.
Validate before expose. We notice that exposing dirty data must be done carefully, especially when combined with no detect actions, as aborting a transaction that has exposed its data can result in a long chain of cascading aborts. Thus, we follow existing work (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/eurosys/MuAS19/DRP, ; DBLP:conf/sosp/XieSLAK015/Callas, ) to perform an early validation before exposing dirty data.
Learn transaction backoff. We follow existing studies (DBLP:conf/sigmod/ShengTZP19/abort_prediction, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ) to learn a backoff for transaction abort and retry. This number is learned together with timeout, for each type of transaction.
4.4. Function Update
The agent function is updated by optimizing a new function to replace the existing one when workload drifts are detected. CCaaLF employs a simple, yet effective threshold-based method for workload drift detection. Specifically, it monitors system performance during database operations and triggers the optimization process upon detecting significant performance changes. The threshold for detection is 10% (in throughput). This value is determined empirically to ensure robust detection and to avoid over-triggering. We will show that our detection method is effective in Section 6.2. We emphasize that CCaaLF only focuses on the design of the concurrency control algorithm. Thus, the problem of workload drift detection is orthogonal. In other words, a detailed analysis of how our detection method compares to other algorithms, such as those discussed in  (DBLP:journals/pacmmod/WuI24/2sample, ), is beyond the scope of this paper.
CCaaLF enables agent function updates without blocking the ongoing or incoming transactions. In particular, each transaction is assigned a specific agent function, namely the latest one, at the start of its execution. During an agent function update, CCaaLF maintains both the latest function and the previously loaded functions still in use by ongoing transactions. While a new function is being loaded, the old agent function continues to handle transactions. Once the loading is complete, the system performs an atomic pointer update to switch to the new agent function. This design avoids blocking ongoing or incoming transactions. It also ensures agent function consistency, that is, it prevents the switch of functions in the middle of a transaction’s execution.
4.5. Detailed Algorithm
We describe how CCaaLF executes an operation based on the learned actions. Given an operation $op$ of transaction $T$, the algorithm first determines the CC actions using the learned function implemented by the GetCC function defined below.
(2) $GetCC(T,op)$ $\mapsto\mathbf{{a}}=\langle\mathbf{{a}}_{d},\mathbf{{a}}_{t},\mathbf{{a}}_{p},(w_{1},\dots,w_{n},\textit{expose})\rangle$
(3) where $\mathbf{{a}}_{d}\in\{\text{{no detection}},\text{{detect critical}},\text{{detect all}}\}$
(4) $\mathbf{{a}}_{t}\in[0,+\infty),\quad\mathbf{{a}}_{p}\in[0,1]$
(5) $w_{i}\in\mathbb{N},\quad expose\in\{0,1\}$
$\mathbf{{a}}_{d}$, $\mathbf{{a}}_{t}$, and $\mathbf{{a}}_{p}$ are the conflict detection method, timeout and transaction priority, respectively. $(w_{1},\dots,w_{n})$ represents the rule for identifying critical conflicts (as explained in Section 2.1). Specifically, let $n$ be the number of transaction types, $w_{i}$ means that $T$ considers the first $w_{i}$ operations of a dependent transaction of type $i$ as critical conflicts. The boolean parameter $expose$ determines whether $T$ makes its operations visible to other transactions after executing $op$.
1 Function Execute($T,op$):
2       a $\leftarrow$ GetCC($T,op$)
3       $op.priority\leftarrow$ ap , $op.tx\leftarrow T$
4
5      if ad = no detection  then
6             $op.clean=True$, $S\leftarrow\varnothing$
7
8      else if ad = detect all  then
9             $S\leftarrow$ all operations from conflict transactions
10      else if ad = detect critical  then
11             if IsStoredProcedure() then
12                   $S\leftarrow\varnothing$ // Conflicts arising from pipeline waits.
13                   for $T^{\prime}\in T.dep$  do
14                        bar $\leftarrow$ $w_{T^{\prime}.type}$ $S\leftarrow S\cup\{\text{1-st}\text{ to }\text{bar-th}\text{ operations in }T^{\prime}\}$
15
16            else if  EarlyValidation$(T)$ = Abort  then
17                  return Abort, $\varnothing$
18
19
20      $S\leftarrow S-\{op^{\prime}\mid op^{\prime}\in S\textup{ {and} }op^{\prime}.priority<op.priority\}$
21       if Not all $op^{\prime}\in S$ finish within timeout at  then
22            return Abort, $\varnothing$
23      $result\leftarrow$ SafeExecute($T,op$)
24
25      if IsStoredProcedure() and $expose=1$  then
26             if  EarlyValidation$(T)$ = Abort  then
27                  return Abort, $\varnothing$
28            else
29                   if $op$ is not the last operation in $T$  then
30                         $\mathbf{a}^{\prime}\leftarrow$ the actions for the next operation in $T$
31                         Perform the pipeline wait actions in $\mathbf{a}^{\prime}$
32
33                  else
34                         Wait for all $T^{\prime}\in T.dep$ to finish
35
36                   for $op^{\prime}\in T.ws$ do Lock tuple($op^{\prime}.key$)
37                   for $op^{\prime}\in$ dirty reads on tuple($op.key$) do
38                         $T.dep\leftarrow T.dep\cup\{op^{\prime}.tx\}$
39                  for $op^{\prime}\in T.ws$ do
40                         Append $op^{\prime}$ to the version chain as dirty data
41                         Unlock tuple($op^{\prime}.key$)
42
43
44      return Succeed, result
45
46
47 Function SafeExecute($T,op$):
48       if $op.type=$ read then
49             if $op.clean$ then
50                   $op^{\prime}\leftarrow$ latest committed data on tuple($op.key$)
51
52            else
53                   $op^{\prime}\leftarrow$ latest dirty data on tuple($op.key$)
54                   $T.dep\leftarrow T.dep\cup\{op^{\prime}.tx\}$
55            $T.rs\leftarrow T.rs\cup\{op^{\prime}\}$
56             $T.validate\leftarrow T.validate\cup\{op^{\prime}\}$
57             return $op^{\prime}$
58      else
59             $T.ws\leftarrow T.ws\cup\{op\}$
60             return $\varnothing$
61
Algorithm 1 Execute an operation with learned actions
4.5.1. Execution
CCaaLF does not guarantee all conflicts are detected and resolved during execution, similar to OCC. A transaction $T$ has the following fields:
  • $T.rw$ and $T.ws$: the local read and write sets.
  • $T.dep$: the set of transactions that $T$ depends on.
  • $T.validate$: operations of $T$ that have been executed but not yet validated.
For each operation $op$, we maintain the following information:
  • $op.tx$: the transaction that contains $op$.
  • $op.key$: the primary key of the data object accessed by $op$.
  • $op.type$: the operation type (read/write).
  • $op.clean$: whether $op$ reads the latest committed (clean) data.
  • $op.priority$: the priority of $op$ when resolving its conflicts with others operations.
Algorithm 1 describes how a transaction operation is executed with learned actions. It performs early validation to proactively abort transactions that will fail later (line  1). The algorithm also enforces early validation before exposing a write (line 1), where the failure of a transaction $T$ could lead to the cascading abort of all transactions depending on it. By performing early validation (Algorithm 2), it reduces the impact of cascading aborts. The algorithm does not perform early validation for “detect all” because doing so incurs additional overhead and can restrict the search space. For example, performing early validation in “detect all” would prevent CCaaLF from replicating the behavior of 2PL, which operates without the overhead of early validation checks. Furthermore, our algorithm distinguishes between stored procedure mode and interactive transaction mode. In the former, all the operations of a transaction are known prior to transaction execution. In the latter, they are not known before transaction execution.
For interactive transactions, the algorithm disallows certain actions that could compromise correctness due to the transaction’s logic. Such actions include dirty reads and partial retry, because the behavior of the conflicting transaction is unknown, and because the user has already seen the outputs of some operations. Before executing $op$, the algorithm calls $GetCC(T,op)$ to get actions $\mathbf{{a}}$ (line 1). Next, it executes conflict detection and stores the conflicts in $S$. This step is skipped if ${\mathbf{{a}}}_{d}$ is no detection, and the algorithm enforces $op$ to read clean data (lines 11). If ${\mathbf{{a}}}_{d}$ is detect critical, it performs early validation to check if previously read data have changed (line 1) and aborts $T$ if the validation fails. The conflicts with lower priorities than that of $op$ are removed from $S$. The remaining are resolved with the given timeout ${\mathbf{{a}}}_{t}$ (lines 11). If not all the conflicts in $S$ are resolved within ${\mathbf{{a}}}_{t}$, $T$ is aborted (line 1). Otherwise, $T$ proceeds to read the latest committed data (line 1), or to write to its write set $T.ws$ (line 1).
For stored procedure transactions, the algorithm enables more flexibility by allowing dirty reads. It uses $w_{1},\dots,w_{n}$ for detecting critical conflicts before dirty reads, and $expose$ for deciding whether to make uncommitted write visible to other transactions. Each data object has a version chain containing dirty data. The execution when ${\mathbf{{a}}}_{d}$ is no detection or detect all is the same as that of interactive transactions. When ${\mathbf{{a}}}_{d}$ is detect critical, it detects critical conflicts between $T$ and its dependent transactions in $T.dep$ (lines  11). In particular, for every $T^{\prime}\in T.dep$, it adds the first $w_{T^{\prime}.type}$ operations of $T^{\prime}$ to $S$. Once all the conflicts in $S$ are resolved, it lets $op$ read the data at the tail of the corresponding version chain, and updates $T.dep$ to track the read-write dependency (lines 11). After executing $op$ (line 1), if $expose=1$, it performs early validation and aborts $T$ if the validation fails. If the validation succeeds, it exposes the uncommitted data in $T.ws$ by appending them to the corresponding version chains.
1 Function EarlyValidation($T$):
2       for $op\in T.validate$  do
3             $op^{\prime}\leftarrow$ latest data version on tuple($op.key$).
4             if $op.tx\neq op^{\prime}.tx$ then
5                  return Abort
6
7      $T.validate=\varnothing$
8       return Succeed
9 Function Commit($T$):
10       for $T^{\prime}\in T.dep$  do
11             Wait $T^{\prime}$ to finish.
12             if $T$ read dirty from $T^{\prime}$ and $T^{\prime}$ gets aborted then
13                  return Abort
14
15      for $op\in T.ws$  do  Lock tuple($op.key$).
       *// Serialization point.
16       for $op\in T.rs$  do
17             $op^{\prime}\leftarrow$ latest committed data version on tuple($op.key$)
18             if $op.tx\neq op^{\prime}.tx$ then
19                   return Abort
20
21      for $op\in T.ws$  do
22             Clean up $T$’s dirty data on tuple($op.key$)
23             Append $op$ to tuple($op.key$)’s version chain as clean data
24             Unlock tuple($op.key$).
25      return Succeed
Algorithm 2 Validation and commit
4.5.2. Validation and commit
Algorithm 2 shows how the algorithm performs an OCC-like four-step validation at commit time. First, it waits for all dependent transactions in $T.dep$ to complete, and aborts if any dependent transaction aborts (lines 22). Second, it locks the private write set to avoid concurrent read of the data in the set (line 2). Third, it verifies that data in the read set are the same as the latest committed versions (lines 22). It aborts if the verification fails. Fourth, it removes dirty data from $T$ on the corresponding version chains and unlocks the write set (lines 22). The latest committed data are updated with those in the write set.
4.6. Proof of Correctness
To prove that CCaaLF ensures serializability, we show that any transaction schedule allowed by CCaaLF can be transformed into a serial schedule in which the committed transactions are executed one by one in isolation. In particular, we show that the transaction execution results are the same when running with the CCaaLF schedule as when running with the serial schedule.
Definitions and Setup:. Let:
  • $T_{1},T_{2},\dots,T_{n}$ be a set of transactions in the system.
  • A schedule $S$ is an ordered sequence of read/write operations from these transactions.
  • A schedule is serial if all operations of a transaction $T_{i}$ execute without interleaving with operations from other transactions.
  • A schedule is serializable if it is equivalent to some serial schedule, i.e., it produces the same final state of the system.
We aim to prove that the schedule $S$ produced by CCaaLF is serializable, i.e., there exists a serial schedule $S^{\prime}$ that is equivalent to $S$.
Transformation to Serial Schedule:.
  1. (1)
    Serialization Point: Denote the serialization point of a transaction $T_{i}$ by $t(T_{i})$. This is the logical time at which $T_{i}$ is considered to have committed in the global order of transactions. This point corresponds to the time when $T_{i}$ is validated in the algorithm (Algorithm 2, line 2).
  2. (2)
    Correctness Condition: We need to show that all operations from the schedule $S$ can be transformed into a serial schedule $S^{\prime}$, without changing the results of the transaction execution. We do this by shifting operations to their serialization points.
  3. (3)
    Read Operations: Let $r(x)$ denote a read operation on data item $x$ by transaction $T_{i}$.
    • Since read operations do not modify data, moving $r(x)$ to the serialization point $t(T_{i})$ does not change the final state of any data objects.
    • Moreover, the validation step (lines 22) ensures that the value read by $r(x)$ is consistent. Specifically, it guarantees that $x$ was not modified between the read and the serialization point.
    • Hence, moving $r(x)$ to $t(T_{i})$ preserves the correctness of the read.
  4. (4)
    Write Operations: Let $w(x)$ denote a write operation on data item $x$ by transaction $T_{i}$.
    • Write operations are protected by locks (lines 22). Therefore, their order is preserved with respect to other conflicting writes, ensuring that moving $w(x)$ to $t(T_{i})$ does not change the final state of any data objects.
    • Since read operations are validated against the committed state (ensured by lines 22), no read operation $r(x)$ from $T_{j}$ can access a value written by a transaction $T_{i}$ unless $T_{i}$’s serialization point precedes the validation of $T_{j}$.
    • In this case, the successful validation also indicates that no other write on $x$ exists in $S$ from $T_{i}$’s serialization point $t(T_{i})$ to $T_{j}$’s $r(x)$.
    • Therefore, moving $w(x)$ to $t(T_{i})$ does not affect any reads in $S$.
Since we can shift all operations in $S$ to their respective serialization points without altering the execution results, the final state of the system under $S$ is identical to that of a serial schedule $S^{\prime}$, where transactions are executed one-by-one according to their serialization points. Thus, every schedule $S$ allowed by CCaaLF is serializable. This completes the proof that CCaaLF ensures serializability.
5. Learning Agent Function
In this section, we first define the optimization of the agent function as a black-box optimization problem. We then describe how CCaaLF efficiently searches for the optimal agent functions.
5.1. Problem Definition
CCaaLF starts the optimization process for the agent function when it detects performance degradation caused by workload drifts. It uses throughput as the performance metrics, similar to previous works (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/cidr/TangJE17/ACC, ).
Let $\mathbf{S}$ be the set of possible input states $\mathbf{{s}}$, and $\mathbf{A}$ be the set of possible output actions $\mathbf{{a}}$. The optimization problem is defined as:
(6) maximize $\text{Score}(\mathcal{{F}})$
(7) subject to $\forall\mathbf{{s}}\in\mathbf{S},\ \mathbf{{a}}=\mathcal{{F}}(\mathbf{{s}})\in\mathbf{A}.$
We observe that different actions are amenable to different optimization techniques. Model-based optimizations are suitable for some actions, whereas search-based techniques are suitable for others. Furthermore, different actions have different impacts on the overall throughput, and thus the order in which they are optimized can affect convergence speed. Based on these observations, we decompose $\mathcal{{F}}$ into three sub-functions: $\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, and $\mathcal{{F}}_{P}$, and optimize them separately. Their outputs are combined to obtain $\mathbf{{a}}$ $=\langle{\mathbf{{a}}}_{d},{\mathbf{{a}}}_{t},{\mathbf{{a}}}_{p},({w_{1}},\dots,{w_{n}},{expose})\rangle$. $\mathcal{{F}}_{T}$ controls the timeout values, corresponding to ${\mathbf{{a}}}_{t}$. $\mathcal{{F}}_{D}$ controls the conflict detection and priority, corresponding to (${\mathbf{{a}}}_{d},{\mathbf{{a}}}_{p}$). $\mathcal{{F}}_{P}$ controls pipeline-wait actions, corresponding to $({w_{1}},\dots,{w_{n}},{expose})$.
Learning the optimal agent function requires evaluating the value of $\mathcal{{F}}$). Existing works (DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/cidr/TangJE17/ACC, ) load the function into the running database and record the system throughput as the performance scores. This approach is time-consuming. In particular, the process of loading and running the system to evaluate $\mathcal{{F}}$) takes seconds  [^1] This process takes 3.68 seconds on average under our default experimental setting, as detailed in Section 6.1: TPCC workload, 16 threads, one warehouse. to complete and is repeated at every optimization step, adding significant overheads to the learning process (DBLP:conf/osdi/WangDWCW0021/polyjuice, ). In addition, the function $\mathcal{{F}}$) is non-smooth due to its discrete input $\mathcal{{F}}_{P}$. This non-smoothness makes it difficult to optimize $\mathcal{{F}}_{P}$ with model-based optimization methods. The following subsections discuss how CCaaLF addresses the aforementioned limitations of the existing approaches for evaluating $\mathcal{{F}}$) in two ways. First, it employs an online-trained surrogate model to minimize the number of performance score evaluations (Section 5.2). Second, it uses a graph reduction algorithm to effectively optimize $\mathcal{{F}}_{P}$ (Section 5.3).
5.2. Agent Function Learning with Online-Trained Surrogate Model
As $\mathcal{{F}}$) is costly to evaluate, CCaaLF treats it as a black-box (regis2013combining/DYCORS, ; DBLP:conf/kdd/GolovinSMKKS17/Vizier, ; DBLP:books/daglib/0022247, ) and adopts a Bayesian optimization approach. The Bayesian optimization can efficiently identify high-performing agent functions with minimal evaluations, thereby reducing the evaluation cost. In particular, CCaaLF employs a probabilistic surrogate model, i.e., the Gaussian process model, which is trained online to approximate $\mathcal{{F}}$). The Gaussian process model is differentiable for efficient optimization and is well-suited as a surrogate model because it predicts both the mean and variance of performance scores, balancing exploration of new areas in the function space with exploitation of known high-performing areas.
This surrogate model guides the optimization process by identifying the most promising agent function, i.e., the one expected to yield the highest performance score improvement, for subsequent deployment and evaluation. Specifically, during an optimization step, the surrogate model is first fitted to the historical data collected since the last workload drift, which optimizes its mean vector and covariance matrix to maximize the data likelihood. Next, for an unevaluated agent function $\mathcal{{F}}$, the fitted model predicts the mean $\mu(\text{Score}(\mathcal{{F}}))$ and variance $\sigma(\text{Score}(\mathcal{{F}}))$ of its performance score. After that, the Upper Confidence Bound (UCB) acquisition function computes a score for $\mathcal{{F}}$ as $\text{UCB}(\mathcal{{F}})=\mu(\text{Score}(\mathcal{{F}}))+\lambda\sigma(\text{Score}(\mathcal{{F}}))$, where $\lambda$ is a trade-off parameter and is set to $2.576$ in experiments, corresponding to the 99th percentile of a standard normal distribution to achieve an effective balance between exploration and exploitation. Note that the UCB score is differentiable with respect to the parameters of $\mathcal{{F}}$, and therefore CCaaLF can explicitly optimize these parameters to maximize the UCB score using the L-BFGS-B (DBLP:journals/mp/LiuN89, ) algorithm, which produces $\mathcal{{F}}_{next}$, the agent function with the highest expected performance score. Once $\mathcal{{F}}_{next}$ is identified, it is evaluated in the database system, and its actual performance score is recorded. Finally, the new data point of $\mathcal{{F}}_{next}$ and its score are added to the historical data, and the surrogate model is updated incrementally. This process is repeated iteratively, progressively refining predictions of $\mathcal{{F}}$) and identifying better agent functions over time.
5.3. Pipeline-Wait Action Optimization via Learned Conflict Graph
The Bayesian optimization technique in Section 5.2 can optimize $\mathcal{{F}}_{T}$ and $\mathcal{{F}}_{D}$. However, optimizing $\mathcal{{F}}_{P}$ is more challenging, especially under complex workloads, such as TPC-C, which involves many potential dependency cycles. Through extensive empirical analysis, we observe that fixing $\mathcal{{F}}_{T}$ and $\mathcal{{F}}_{D}$ while optimizing $\mathcal{{F}}_{P}$ leads to poor locality of the black-box function Score($\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, $\mathcal{{F}}_{P}$) around the optimal solutions, i.e., a small change to $\mathcal{{F}}_{P}$ can lead to significant performance reduction due to new dependency cycles. Pipeline-wait actions, which are parameterized by $(w_{1},\dots,w_{n},expose)$, are designed to avoid such cycles because they lead to cascading aborts or deadlocks (DBLP:conf/sigmod/WangMCYCL16/ic3, ). As a result, the surrogate model trained via Bayesian optimization fails to approximate Score($\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, $\mathcal{{F}}_{P}$) accurately, because the optimization inherently assumes smoothness and locality in the approximated function. Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) discusses similar issues with Policy Gradient optimization which motivates it to adopt a genetic search algorithm.
1
2 Function GraphReduction($branchFactor,mutateRate,K$):
3       $\mathcal{{V}}_{\textup{false}}\leftarrow$ a default function that returns false for all inputs
4       ${\mathcal{{F}}}_{best}\leftarrow$ the current best agent function.
5       $Pop\leftarrow\{(\mathcal{{V}}_{\textup{false}},\mathcal{{V}}_{\textup{false}},initialScore)\}$
6       while $Pop\neq\varnothing$  do
7             for $(\mathcal{{V}}_{e},\mathcal{{V}}_{o},score)\in Pop$ do
8                   for $i\leftarrow 1$ to branchFactor  do
9                         $(\mathcal{{V}}_{e}^{\prime},\mathcal{{V}}_{o}^{\prime},score^{\prime})\leftarrow\textup{Mutate(}\mathcal{{V}}_{e},\mathcal{{V}}_{o},mutateRate$)
10                         $nTry\leftarrow 1$
11                         while ($\mathcal{{V}}_{e}^{\prime},\mathcal{{V}}_{o}^{\prime}$) has been searched before and $nTry\leq MaxTry$ do
12                               $(\mathcal{{V}}_{e}^{\prime},\mathcal{{V}}_{o}^{\prime},score^{\prime})\leftarrow\textup{Mutate}(\mathcal{{V}}_{e},\mathcal{{V}}_{o},mutateRate)$
13                               $nTry\leftarrow nTry+1$
14                        if $nTry>MaxTry$ then
15                               $Pop\leftarrow Pop-\{(\mathcal{{V}}_{e},\mathcal{{V}}_{o},score)\}$
16                               break
17                        else
18                               $Pop\leftarrow Pop\cup\{(\mathcal{{V}}_{e}^{\prime},\mathcal{{V}}_{o}^{\prime},score^{\prime})\}$
19
20
21
22            $Pop\leftarrow$ GetTopK($Pop$)
23             if $Pop$ did not change in $MaxStep$ steps  then
24                   break
25
26      return ${\mathcal{{F}}}_{best}$
27
28
29 Function Mutate($\mathcal{{V}}_{e},\mathcal{{V}}_{o},prob$):
       // Mutate by merging nodes.
30       for $((T,op)\to expose)\in\mathcal{{V}}_{e}$  do
31             Mutate $expose$ to true with $prob$ probability
32
      // Mutate by cutting edges.
33       for $((T,op)\to optimistic)\in\mathcal{{V}}_{o}$ do
34             Mutate $optimistic$ to true with $prob$ probability
35
36      ${\mathcal{{F}}}_{new}\leftarrow\text{UpdateWaitActions}(\mathcal{{V}}_{e},\mathcal{{V}}_{o},{\mathcal{{F}}}_{best})$
37       if $\textup{Score}({\mathcal{{F}}}_{new})>$ Score(${\mathcal{{F}}}_{best}$)  then
38             ${\mathcal{{F}}}_{best}\leftarrow{\mathcal{{F}}}_{new}$
39
40      return $\mathcal{{V}}_{e},\mathcal{{V}}_{o},\text{Score}({\mathcal{{F}}}_{new})$
41
42
Algorithm 3 Graph reduction search algorithm
To address the challenge of global optimization of $\mathcal{{F}}_{P}$, CCaaLF learns an optimal conflict graph from which the pipeline-wait actions can be derived, instead of learning these actions directly. This reformulation simplifies the learning process and improves optimization performance. A conflict graph is a static graph generated from the workload. In this graph, each type of transaction operation (e.g., a line of SQL) is a node. If two operations have potential conflicts (e.g., read-write conflicts), an edge is added between the corresponding nodes. CCaaLF uses SC-graph, a conflict graph widely adopted in existing CC algorithms (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/osdi/MuCZLL14/Rococo, ). These algorithms build a conflict graph offline and compute pipeline-wait actions by analyzing this graph prior to workload execution. However, they rely on static conflict graphs, which do not account for dynamic factors such as access skewness. We contend that the conflict graph needs to be dynamic, with conflict edges forming based on certain probabilities. For example, in the TPC-C workload, if there are many warehouses and few concurrent users, the update WAREHOUSE table access in the payment transaction is unlikely to conflict with the new order transaction’s read WAREHOUSE table access, and therefore this edge can be deleted. However, when there are few warehouses and many concurrent users, these accesses are more likely to conflict, therefore adding the conflict edge is beneficial.
1
2 Function GetWaitActions($op,T,\mathcal{{V}}_{e},\mathcal{{V}}_{o},LastExposed$):
3       $G_{\text{cut}}\leftarrow\text{CutEdgesAroundOptimisticOperations}(G_{\text{full}},\mathcal{{V}}_{o})$
4       $n\leftarrow$ number of operations in $T$
5       $lastExposePoint\leftarrow n+1$
6       $wait_{op,T.\text{type}}\leftarrow 0$
       // Get wait actions.
7       for $i\leftarrow n$ to 1 do
8             $op^{\prime}\leftarrow$ the $i_{\text{th}}$ operation of $T$
9             if $\mathcal{{V}}_{e}(T,op^{\prime})=\textbf{{false}}$ then
10                  $lastExposePoint\leftarrow i$
11            if $\langle op,op^{\prime}\rangle\in G_{\text{cut}}$ and $op^{\prime}$ is read then
12                   $wait_{op,T.\text{type}}\leftarrow i$
13                   break
14            if $\langle op,op^{\prime}\rangle\in G_{\text{cut}}$ and $op^{\prime}$ is write then
15                   $wait_{op,T.\text{type}}\leftarrow lastExposePoint$
16                   break
17
      // Defer the wait actions of write operations until they are exposed to concurrent transactions.
18       if $op$ is read  then
19             $res\leftarrow wait_{op,T.\text{type}}$
20
21      else
22             $res\leftarrow 0$
23
24      if $LastExposed\neq\varnothing$ then
25             for $op^{\prime}\in LastExposed$ do
26                   if $wait_{op^{\prime},T.\text{type}}>res$ and $op^{\prime}$ is write then
27                         $res\leftarrow wait_{op^{\prime},T.\text{type}}$
28
29
30
31      return $res$
32
Algorithm 4 Pipeline wait actions
CCaaLF employs a graph reduction algorithm inspired by genetic search to quickly learn the optimal conflict graph. The algorithm starts with a large static conflict graph $G_{full}$, where each node represents a transaction operation $op$ from a transaction $T$, and an edge connecting two operations accessing the same table and least one operation is a write. The modifications of the conflict graph $G_{full}$ are defined by two functions, $\mathcal{{V}}_{e}$ and $\mathcal{{V}}_{o}$:
  • $\mathcal{{V}}_{e}(T,op)=$ true if the node corresponding to operation $op$ in $T$ is merged with the node corresponding to the operation next to $op$ in $T$.
  • $\mathcal{{V}}_{o}(T,op)=$ true if edges involving the node corresponding to $op$ in $T$ are removed.
1
2 Function UpdateWaitActions($\mathcal{{V}}_{e},\mathcal{{V}}_{o},{\mathcal{{F}}}_{best}$):
       // Get pipeline wait actions for all transaction operations.
3       for $T\in$ all types of transactions do
4             $n\leftarrow$ number of operations in $T$
5             for $T^{\prime}\in$ all types of transactions do
6                   $unexposed\leftarrow\varnothing$
7                   $toExpose\leftarrow\varnothing$
8                   for $i\leftarrow 1$ to $n$ do
9                         $op\leftarrow$ the $i_{\text{th}}$ operation of $T$
10                         $WaitAct_{T,T^{\prime},i}\leftarrow$ GetWaitActions($op,T^{\prime},\mathcal{{V}}_{e},\mathcal{{V}}_{o},toExpose$)
11                         $unexposed=unexposed\cup\{op\}$
12                         $toExpose=\varnothing$
13                         if $\mathcal{{V}}_{e}(T,op)=$ false then
14                               $toExpose=unexposed$
15                               $unexposed\leftarrow\varnothing$
16
17
18
19
20      $\mathbf{S}\leftarrow$ GetInputSet(${\mathcal{{F}}}_{best}$)
21       $\mathcal{{F}}_{new}\leftarrow{\mathcal{{F}}}_{best}$
       // Update the pipeline wait actions in $\mathcal{{F}}_{new}$.
22       for $s\in\mathbf{S}$ do
23             $isOptimistic\leftarrow$ true
24             for $T^{\prime}\in$ all types of transactions do
25                   $\mathcal{{F}}_{new}(s).w_{T^{\prime}.type}\leftarrow 0$
26                   for $T$ whose $i_{th}$ operation can produce $s$ state do
27                         $op\leftarrow$ the $i_{\text{th}}$ operation of $T$
28                         if $WaitAct_{T,T^{\prime},i}>\mathcal{{F}}_{new}(s).w_{T^{\prime}.type}$  then
29                               $\mathcal{{F}}_{new}(s).w_{T^{\prime}.type}\leftarrow WaitAct_{T,T^{\prime},i}$
30
31                        if $\mathcal{{V}}_{o}(T,op)=$ false then
32                               $isOptimistic\leftarrow$ false
33
34
35
36            if $isOptimistic=$ true  then
37                   $\mathcal{{F}}_{new}(s).\mathbf{a}_{d}\leftarrow$ no detect
38
39      return ${\mathcal{{F}}}_{new}$
40
Algorithm 5 Updating pipeline wait actions
As outlined in Algorithm 3, the graph search algorithm starts with no conflict graph modifications, $\mathcal{{V}}_{e}=\mathcal{{V}}_{o}=\mathcal{{V}}_{\text{false}}$, then searches for better conflict graphs through mutations (lines 33 and lines 33). The algorithm maintains a population of the top $K$ best conflict graphs found during the search. The conflict graphs are evaluated by converting them into pipeline-wait actions (Algorithm 4), generating new agent functions with these pipeline-wait actions (Algorithm 5), and evaluating the corresponding agent functions. Specifically, given a conflict graph, we identify all potential dependency cycles and compute the required pipeline-wait actions to resolve these cycles (Algorithm 4 lines 44). After this, we update the pipeline-wait actions in the current best agent function to our calculated pipeline-wait actions and evaluate its performance score as the score for the conflict graph (Algorithm 3 line 3). As an agent function input may map to multiple nodes in the conflict graph, we select the longest wait action to prevent potential dependency cycles (Algorithm 5 lines 55). We adopt the same optimizations as IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, ) to enforce the pipeline wait actions of the next operation before exposing dirty writes (Algorithm 1 lines 11), and to defer the pipeline waits of unexposed write operations until they are exposed (Algorithm 4 lines 44).
During optimization, the algorithm iteratively improves the conflict graphs by mutating the current graph population to better capture actual conflicts. Each mutation simplifies the graph by merging two consecutive nodes into a single node, or by removing conflict edges (Figure 6 and Algorithm 3). The mutation aims to eliminate unnecessary conflicts that will be resolved during the pipeline-wait actions calculation. Since the mutation reduces the number of nodes or the number of edges, the size of the graph reduces quickly towards the optimal conflict graph. The graph reduction search algorithm stops when no further mutations yield performance improvements (Algorithm 3 line 3), or when all mutations of the current population have been exhausted (line 3).
The graph reduction search algorithm complements the Bayesian optimization for $\mathcal{{F}}_{P}$. The former is efficient at exploring potentially optimal agent functions derived from conflict graphs, which Bayesian optimization alone may struggle to explore. CCaaLF still relies on Bayesian optimization to approximate high-performance but sparsely distributed agent functions. For instance, in the YCSB workload, we observe that a no-pipeline-wait policy can outperform learned pipeline-wait actions due to reduced computational costs associated with conflict detection and dependency tracking. Such an agent function cannot be identified through graph reduction search, but it can be found quickly by Bayesian optimization within minutes.
Figure 6. Two types of graph reduction mutations.
5.4. Agent Function Optimization Pipeline
Given a time budget, it is important to prioritize the optimization of actions that are more likely to yield significant performance improvements. CCaaLF uses an optimization pipeline that transfers knowledge from one optimization stage to the next. It allocates the time budget efficiently across these stages. Recall that the agent function $\mathcal{{F}}$ is decomposed into three sub-functions $\mathcal{{F}}_{P}$, $\mathcal{{F}}_{D}$, and $\mathcal{{F}}_{T}$ (see Section 5.1). The optimization pipeline consists of four stages, each targeting a subset of actions while keeping the remaining actions unchanged from the best-seen agent function. Each stage leverages the previously evaluated agent functions and their corresponding performance scores as initial training data for the stage’s optimizer.
The first stage optimizes both $\mathcal{{F}}_{P}$ and $\mathcal{{F}}_{D}$ using the graph reduction algorithm. These two sub-functions are optimized first because they lead to the largest performance gains, e.g., with improvements of nearly 80% in the TPC-C workload. $\mathcal{{F}}_{P}$ and $\mathcal{{F}}_{D}$ can be optimized simultaneously, as actions no detection and cut-edge mutations both assume no conflicts around an operation. This stage starts with the same initial configuration as that of IC3, in which all conflict detection actions are set to detect critical. The graph reduction search iteratively refines the conflict graph. When an agent function input $s$ maps to conflict graph nodes that are entirely disconnected from others, it indicates that database operations in state $s$ rarely encounter conflicts. In such cases, we set the conflict detection action to no detect to reduce transaction execution costs (Algorithm 5 lines 55). The graph population size is set to $K=4$, enabling quick convergence to an efficient agent function. The stage ends when no further mutations yield performance improvements, typically within 10 minutes due to the monotonically reducing nature of the mutation process.
The second stage optimizes $\mathcal{{F}}_{T}$ using Bayesian optimization. By this stage, most parameters in $\mathcal{{F}}_{T}$ are continuous and exhibit good locality. This stage ends when no further performance improvements are observed after several consecutive evaluations (e.g., 20 evaluations). The third stage is the same as the first stage, but with a larger population size ($K=8$). This stage allows for deeper exploration of the conflict graph, in order to refine the conflict graph and agent functions. Finally, the fourth stage fine-tunes all the parameters of $\mathcal{{F}}$ using Bayesian optimization. This stage continues until the time budget is exhausted.
5.5. Feaure Selector Optimization
We formulate the feature selector as a function that maps a feature selection strategy to the performance score achieved by CCaaLF with the corresponding selected features. Each feature selection strategy is represented as a vector, with each element indicating whether a corresponding feature is included and specifying its transformation type (linear, square root, or logarithmic). The output of the feature selection strategy is a set of input features for the agent function $\mathcal{F}$, which is subsequently used during the agent function optimization. Similar to the agent function optimization described in Section 5.2, Bayesian optimization is employed to iteratively propose new feature selection strategies, evaluate them, and refine the surrogate model. This process identifies the set of features that maximizes the performance score.
To balance effectiveness and efficiency, the feature selection strategy is optimized to maximize the performance improvement achieved through function optimization within a user-defined time budget. Note that selecting too few features can overly constrain the search space of agent functions, leading to suboptimal solutions, whereas selecting too many features can expand the search space excessively, making the optimization computationally expensive. Therefore, we balance the two factors by incorporating the time budget into the performance score. We allocate a default budget of one hour and run the feature selector optimization for one day to determine the feature selection strategies used in our experiments. Because the optimization is costly, we pre-optimize a selector that performs well under the default settings. The user can opt to run the optimization to further improve the default selector’s performance. The optimized feature selectors and their configurations are available in our code repository (CCaaLF-impl, ).
6. Performance Evaluation
In this section, we evaluate the performance of CCaaLF and compare it with state-of-the-art concurrency control (CC) algorithms under diverse workloads. We implement CCaaLF in C++ and integrate it into Silo (DBLP:conf/sosp/TuZKLM13/silo, ), a widely used database system for evaluating CC algorithms. Specifically, we add the following components to the database engine: a collector that captures transaction-related features, a feature selector that refines the captured features, an in-database cache that stores the learned agent functions, and a monitor that tracks system performance for function optimization and workload drift detection. We modify the concurrency control code in the database engine to integrate the logic presented in Algorithm 1. We further implement a script that processes optimization requests and training data from the database. Our code is available at (CCaaLF-impl, ).
6.1. Experimental Setup
The experiments are conducted on a 24-core Intel(R) Xeon(R) Gold 5317 server with 123GB of memory. The CPU runs at a base frequency of 800 MHz, and a maximum turbo frequency of 3.6 GHz.
Execution. CCaaLF executes transactions in one of the following two modes: stored procedure mode (DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ; DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/sigmod/GuoWYY21/bamboo, ; DBLP:conf/sigmod/LinC0OTW16/LEAP, ; DBLP:journals/vldb/YaoZLOX18/DGCC, ) and interactive transaction mode (DBLP:journals/tocs/ZhangSSKP18/TAPIR, ; DBLP:journals/tocs/CorbettDEFFFGGHHHKKLLMMNQRRSSTWW13/spanner, ; DBLP:conf/sigmod/PavloCZ12/skew_aware, ; DBLP:journals/pvldb/YuBPDS14/1000_core_cc, ).
  • Stored procedure mode. This mode represents how stored procedure transactions are executed by the database engine. In particular, all the transaction’s operations are known before execution. Therefore, concurrency control algorithms can use features such as transaction type and access ID. The transaction only returns results to the user after the transaction commits or aborts, which means it can read uncommitted data during execution and employ database-controlled partial retries.
  • Interactive transaction mode. This mode represents how interactive, user-driven transactions are executed. In particular, the transaction’s operations (or SQL queries) are issued and executed individually. The user waits for the response of each operation before issuing the next. Because the operations are not known, concurrency control algorithms cannot rely on features such as transaction type and access ID. Furthermore, since the results are returned to the user immediately after each operation, CC algorithms cannot perform actions such as partial retry and reading dirty data from uncommitted transactions.
Workloads. We conduct experiments on the widely used YCSB, TPC-C, and TPC-E workloads. We change three types of workload settings for evaluation: access pattern, contention level, and system load. The workloads and their settings are described below.
  • YCSB-extended. We extend YCSB to group multiple read/write operations into a single transaction, and simulate varying access patterns by changing hotspot positions. Specifically, we fix the list of read and write operations for all transactions, representing stable operations found in real-world applications  (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ). We then vary the positions of hotspots, where hotspot accesses follow a highly skewed distribution (Zipfian factor $\theta=1$), while other accesses follow a uniform distribution across all keys (Zipfian factor $\theta=0$). For this workload, we use 16 threads.
  • TPC-C. We use the standard TPC-C workload, which consists of five transaction types: two read-only transactions and three read-write transactions. This workload has complex inter-transaction dependencies, making it harder to optimize than YCSB-extended. We vary the contention level by changing the number of warehouses (fewer warehouses mean higher contention). We vary the system load by changing the number of threads (more threads mean a higher system load). For this workload, the default setting is 16 threads and one warehouse. For a fair comparison, we follow Polyjuice to include all five types of TPC-C transactions. We learn the concurrency control for three types, i.e., payment, new order, and delivery, while excluding the remaining two types, as they are read-only transactions.
  • TPC-E. We employ a TPC-E variant used in (DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ), which includes three read-write transactions trade order, trade update, and market feed. Compared to the TPC-C and YCSB-extended workloads, this workload has more operations per transaction, resulting in a larger search space for learning-based algorithms. We vary the contention level by changing the access skewness, which is controlled via the update operations to the SECURITY table (DBLP:conf/osdi/WangDWCW0021/polyjuice, ), following a Zipfian distribution. The change of the system load setting in the TPC-E workload is the same as that of the TPC-C workload.
We note that both CCaaLF and other learned concurrency control algorithms adapt to workload drifts by first detecting the drifts and then reoptimizing their algorithms. Specifically, the agent functions (in CCaaLF) or policies (in Polyjuice) are retrained when workload drifts are detected.
Baselines. We compare CCaaLF against five state-of-the-art concurrency control algorithms: Silo (DBLP:conf/sosp/TuZKLM13/silo, ), Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ), IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, ), 2PL (2PL_no_wait, ), and CormCC (DBLP:conf/usenix/TangE18/CormCC, ). For the first four baselines, we use the implementations in (DBLP:conf/osdi/WangDWCW0021/polyjuice, ). For CormCC, we follow the approach in Polyjuice and report the best performance between 2PL and Silo.
6.2. Stored Procedure Mode
CCaaLF and all the baselines support the stored procedure mode. We ran all the experiments five times and report the average results. Each algorithm runs for three hours to match the time Polyjuice requires to reach optimal performance. We use IC3 as the initial agent function for agent function optimization.
(a) Access pattern
(b) High contention
(c) Low contention
(d) System load
Figure 7. Performance comparison between CCaaLF and baselines under the YCSB-extended and TPC-C workloads in the stored procedure mode.
Varying access pattern. We fix the contention level and system load while varying the access pattern using four different hotspot distributions. [^2] The values of 1 and 0 indicate hotspot and uniform access, respectively. The four access patterns, from the left to the right in Figure 7(a), are $(0,0,0,1,0,0,0,0,0,0)$, $(1,0,0,1,1,0,0,0,0,0)$, $(0,0,0,1,0,1,0,1,1,0)$, and $(1,1,0,0,1,0,0,1,1,0)$. Figure 7(a) shows the results for the YCSB-extended workload. The hotspot distributions, shown from left to right, shift from being skewed at the beginning of the transaction to being skewed at the end. We observe that CCaaLF has the best throughput, up to 3.32$\times$ higher than that of the second-best baseline (Figure 7(a) Pat. 1). The reason is that CCaaLF learns a set of optimized pipeline-wait actions, timeouts, and conflict detection actions, enabling efficient transaction pipelining for hotspots while reducing the overhead for non-conflicting operations.
It can be seen that the performance gap between CCaaLF and the second-best baseline shrinks as the hotspot distribution shifts toward the end of the transaction. This is because the learned pipeline actions, which contribute most of our performance advantage in this setting, become less effective as the hotspots shift forward. As highlighted in existing works (DBLP:conf/sigmod/GuoWYY21/bamboo, ; DBLP:journals/pvldb/FaleiroAH17/EWV, ), the advantage of dirty reads over clean reads relies on the trade-off between the gain from reducing block time and the overheads from cascading aborts and other issues caused by dirty reads. Therefore, with the hotspots shifting forward, the gain becomes smaller as the transaction is near completion. In other words, the benefit of learned pipeline-wait actions becomes less significant.
Varying contention level. We study the impact of contention levels by fixing the thread number to 16 and setting the number of warehouses to one for high contention (Figure 7(b)) and four for low contention (Figure 7(c)) under the TPC-C workload. We use shaded areas to represent the interquartile range, i.e., the 25th to 75th percentiles of throughput across multiple runs, illustrating the fluctuations of throughput in different CC algorithms. We can see that CCaaLF achieves up to 1.14$\times$ throughput of the second-best baseline. Also, CCaaLF achieves a more stable learning process compared to Polyjuice. Note that Polyjuice relies on random genetic search, which leads to high fluctuations. In contrast, CCaaLF employs an online-trained surrogate model to guide optimization, achieving greater stability throughout the optimization process.
Varying system load. Figure 7(d) shows the results with varying system loads. We fix the number of warehouses at one and increase the thread number from 1 to 16. There are two main observations. First, CCaaLF outperforms all the baselines under high system load (more than four threads). This is because CCaaLF optimizes a better combination of CC actions that allows for higher concurrency. Second, both CCaaLF and Polyjuice perform slightly worse (less than 6.2% drop in throughput) than Silo under low system load (one thread), despite converging to the same CC actions. We note that Polyjuice reported a similar performance drop. The reason is that the learning processes of CCaaLF and Polyjuice incur additional overheads, which do not exist in the non-learned Silo system.
(a) Contention level
(b) System load
Figure 8. Performance comparison between CCaaLF and baselines under the TPC-E workload in the stored procedure mode.
Large benchmark. We further use TPC-E to evaluate CCaaLF against a larger search space. For comparison purposes, we follow  (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) and vary skewness by adjusting the Zipfian distribution of the updates to the SECURITY table. In Figure 8(a), we use 16 threads and vary the Zipfian factor $\theta$ from 0 to 6. As the Zipfian factor $\theta$ increases, accesses to the SECURITY table become more skewed, leading to more conflicts between transactions accessing this table. The results show that CCaaLF consistently outperforms both 2PL and IC3 in these settings. 2PL exhibits suboptimal performance due to the high conflict detection costs inherent in its lock maintenance, whereas IC3 suffers from performance degradation caused by unnecessary pipeline-wait actions derived from its full conflict graph. CCaaLF effectively mitigates both issues by reducing conflict detection costs through learned actions and optimizing the conflict graph to eliminate unnecessary waits.
In Figure 8(b), we fix the Zipfian factor to six and increase the number of threads from 1 to 16. A higher thread number increases the system load and hence the likelihood of transaction conflicts. We observe that while CCaaLF performs worse than Silo in low system load (thread number $\leq$ 8), it outperforms the baselines (up to $\times1$) in high system load (16 threads). As expected, the learning overhead of CCaaLF becomes more significant under low system loads, and becomes less noticeable when the overall performance is greatly improved due to CCaaLF’s ability in identifying efficient CC actions under a high system load.
(a) Workload drift
(b) Impact of threshold value
(c) Impact of features
(d) Impact of learned actions
(e) Time to peak performance
Figure 9. Analysis of CCaaLF in the stored procedure mode.
Workload drift. We evaluate the throughput changes of CCaaLF in response to workload drifts under the TPC-C workload. We simulate workload drifts by changing the number of threads for different time periods, i.e., 1 thread for Setting 1 starting at minute 0, 16 threads for Setting 2 starting at minute 200, and 4 threads for Setting 3 starting at minute 400. Figure 9(a) compares the throughput changes between optimizing from the initial agent function (“Restart”) and optimizing from the previously learned function (“w/o Restart”). We can observe that CCaaLF with “w/o Restart” fails to achieve optimal performance within three hours in both Setting 2 and Setting 3. This is because our graph reduction search algorithm only involves cutting-edge and merging-node mutations. We note that in order to fully explore the conflict graph search space, the optimization must begin with the full conflict graph, containing all nodes and edges. Therefore, starting the optimization from a previously learned optimal agent function, where the conflict graph has already been reduced, restricts the exploration of search space, and leads to a suboptimal conflict graph for the new workload.
Impact of threshold value. We use eight different TPC-C workload settings [^3] Denote $(x,y)$ as $x$ threads and $y$ warehouses, the eight settings are $(1,1)$, $(2,1)$, $(4,1)$, $(8,1)$, $(16,1)$, $(16,2)$, $(16,4)$, and $(16,8)$. and analyze how the threshold value affects the accuracy of drift detection when switching the workload among these settings. We use false positive rate and false negative rate as the accuracy metrics. For the false positive rate, we set the workload to one of the eight settings and report the percentage of experimental runs where a workload drift is falsely detected despite no actual workload drift occurring. For the false negative rate, we evaluate all $7\times 8$ possible workload drifts, starting from one of the eight settings and changing to another, and report the percentage of experimental runs where these drifts are not detected. As shown in Figure 9(b), the threshold value of 10% leads to robust detection (3.6% false negative rate) without over-triggering optimization processes (0% false positive rate).
Impact of features. We conduct an ablation study to evaluate the impact of the two features in the optimized feature set for the TPC-C workload. These features are the transaction type and the number of executed SQLs. Figure 9(c) shows the results under two settings: one warehouse and four warehouses, with features progressively removed from the optimized feature set. It can be seen that both features are necessary for high performance.
Impact of learned actions. We evaluate the impact of the learned actions by reverting the learned CC actions in the final learned agent function to the actions in the initial agent function (IC3). Figure 9(d) shows that different actions have different impacts. Among them, the $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$ sub-functions (see Section 5.1) — ones that are prioritized in our optimization pipeline — account for the majority of the performance improvement (85.8% and 80.6% in the one warehouse and 4 warehouse settings, respectively). The reason is that, although dirty reads allow for a higher degree of concurrency between transactions, they also introduce the risk of dependency cycles that can severely degrade performance. By optimizing $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$, CCaaLF mitigates this risk while ensuring high concurrency.
Optimization time. Figure 9(e) shows the cumulative distribution function of the optimization time for CCaaLF and Polyjuice to reach peak performance. We can observe that CCaaLF has a significantly lower average optimization time (9.67 minutes) compared to Polyjuice (1.78 hours), achieving an 11$\times$ speedup in optimization time. Further, the optimization time of CCaaLF exhibits a lower variance than that of Polyjuice. Unlike Polyjuice which adopts a random genetic search, CCaaLF adopts a graph reduction search guided by an online-trained surrogate model and hence incurs a shorter optimization time with lower variance.
6.3. Interactive Transaction Mode
We compare CCaaLF against three other baselines that support the interactive transaction mode, namely 2PL, Silo, and CormCC. In this mode, we use 2PL as the initial agent function for agent function optimization.
[…]
(a) High contention
(b) Low contention
(c) Read ratio
(d) Transaction length
Figure 10. Performance comparison between CCaaLF and baselines in the interactive transaction mode.
Varying read ratio and transaction length. To understand the impacts of the read ratio and transaction length, we perform two sets of experiments using the YCSB-extended workload. First, we fix the transaction length at 10 and vary the read ratio (i.e., the percentage of read operations) from 0% to 100%. Second, we fix the read ratio at 50% and vary the transaction length from 1 to 16. The results are presented in Figure 10(c) and Figure 10(d), respectively. The throughput of all algorithms increases as the read ratio increases or the transaction length decreases. This trend is expected as higher read ratios or fewer transaction operations result in fewer conflicts among transactions. Further, CCaaLF outperforms the second-best baselines by up to 1.7$\times$ in throughput when the read ratio is below 100%. While under read-only workloads, CCaaLF performs slightly worse (by 2.5%) than Silo. For the same reason as mentioned in Section 6.2, learning-based algorithms can learn complex actions that deliver high performance when there are many conflicts, but the learning overhead becomes significant when there are few conflicts.
(a) Impact of surrogate model
(b) Time to peak performance
Figure 11. Analysis of CCaaLF in the interactive transaction mode.
Impact of surrogate model. We evaluate the impact of the surrogate model on CCaaLF’s throughput by varying the model’s training time under the TPC-C workload. The results for one warehouse and four warehouses are shown in Figure 11(a). As expected, longer surrogate model training times improve the surrogate model’s accuracy. This, in turn, enables CCaaLF to achieve higher performance by identifying better agent functions.
Optimization time. Figure 11(b) depicts the cumulative distribution function of the optimization time for CCaaLF to reach peak performance. Compared to the stored procedure mode (Figure 9(e)), the optimization time in the interactive transaction mode is lower (3.59 minutes on average) because CCaaLF does not need to learn $\mathcal{{F}}_{P}$ for the fact that dirty reads are not permitted in this mode.
(a) Stored procedure mode
(b) Interactive transaction mode
Figure 12. Case studies for learned actions.
6.4. Case Study of Learned Actions
Figure 12 shows two examples that compare the CC actions learned by CCaaLF with those of the second-best baseline, i.e., Polyjuice in the stored procedure mode and 2PL in the interactive transaction mode.
In the stored procedure mode, as shown in Figure 12(a), both CCaaLF and Polyjuice read dirty data during WAREHOUSE table accesses and track the transaction dependency $T_{no}\to T_{pay}\to T^{\prime}_{no}$. Polyjuice learns to perform the STOCK table access of $T^{\prime}_{no}$ only after the STOCK table access of $T_{no}$, preventing a potential dependency cycle $T_{no}\to T_{pay}\to T^{\prime}_{no}\to T_{no}$. In contrast, CCaaLF learns a conflict graph where STOCK table accesses of new order transactions are unlikely to conflict with each other. Therefore, it enables a more efficient transaction interleaving where the STOCK table access from $T_{no}$, the DISTRICT table access from $T_{pay}$, and the STOCK table access from $T^{\prime}_{no}$ can be executed in parallel.
In the interactive transaction mode (Figure 12(b)), 2PL resolves the conflicts between $T_{no}$ and $T_{pay}$ by performing the WAREHOUSE table access of $T_{no}$ only after $T_{pay}$ commits. This results in a long execution time. CCaaLF, on the other hand, learns to order $T_{no}$ before $T_{pay}$ ($T_{no}\to T_{pay}$), allowing $T_{no}$ to optimistically access the WAREHOUSE table without waiting for $T_{pay}$ to commit. It also learns to assign a high priority to the DISTRICT table access of $T_{no}$, ordering this access before that of $T_{pay}$. This prevents a dependency cycle $T_{no}\to T_{pay}\to T_{no}$. The safe interleaving of the two transactions allows them to commit faster in CCaaLF.
7. Related Works
Adaptive CC algorithms are designed to achieve high performance across different workloads. They follow a classify-then-assign workflow: analyzing workload information, classifying transactions based on manually selected features, and assigning the most suitable CC algorithm for each type of transaction. The assignment step can be either heuristic-based (DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/mascots/TaiM96/AEP, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ) or learning-based (DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ).
Among the algorithms adopting heuristic assignment, C3 (DBLP:conf/icde/Su0Z21/C3, ) uses locality-sensitive hashing (LSH) to cluster transactions with similar SQL texts, which tend to conflict on shared data objects. It applies 2PL for intra-group conflicts and OCC for inter-group conflicts. AEP (DBLP:conf/mascots/TaiM96/AEP, ) predicts whether a distributed transaction will encounter conflicts based on local data, then uses EWL when there are no conflicts, and uses PSL when there are. MCC (DBLP:conf/sosp/XieSLAK015/Callas, ) and Tebaldi (DBLP:conf/sigmod/SuCDAX17/Tebaldi, ) classify transactions based on prior workload knowledge, constructing an algorithm hierarchy that switches between RP, SSI, and 2PL, depending on conflict type. ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ) categorizes data into hot, warm, and cold based on the contention level, then uses OCC for cold data, 2PL for hot data, and early validation with retry for warm data. CCaaLF covers more CC designs than these algorithms, and it learns a fine-grained and flexible combination of them for better performance.
CCaaLF has more similarities with the algorithms adopting learning-based assignment. Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) classifies transaction operations by their access ID (a hard-coded unique transaction access identifier) and uses an evolutionary algorithm to optimize pipeline-wait actions and transaction back-offs. ACC (DBLP:conf/cidr/TangJE17/ACC, ) clusters transactions similar to PartCC (DBLP:journals/pvldb/KallmanKNPRZJMSZHA08/HStore, ) and assigns the best algorithm among OCC, 2PL, and PartCC to each cluster. For fast reconfiguration, it also trains a predictive model that predicts the optimal algorithm for each cluster, enabling automatic online reconfiguration with offline-trained models. CCaaLF supports more CC design choices than these algorithms, therefore it achieves higher performance across more diverse workloads. In particular, it does not require prior knowledge of the workloads for access ID assignment or transaction clustering.
8. Conclusions
In this paper, we presented CCaaLF, a novel learned concurrency control algorithm that dynamically adapts to varying workloads by modeling concurrency control as a learnable agent function.
[…]